OpenStudy (anonymous):

For the given function, find the vertical and horizontal asymptote(s) (if there are any). f(x) = (x-5)/(x^2-1) A) x = 5, y = 0 B) x = 1, x = -1, y = 0 C) x = 1 D) x = -1, y = 1

OpenStudy (anonymous):

\[f(x)=\frac{ x-5 }{ x ^{2}-1 }\]

OpenStudy (anonymous):

factorize the denimoninator

OpenStudy (anonymous):

I didn't think you could factor that...

OpenStudy (anonymous):

the denominator....??? u can (x^2-1)=(x+1)(x-1)

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

so, is it b?